### v1

# Cycle products and efficient vectors in reciprocal matrices

Abstract
Given an $n$-by-$n$ reciprocal (pair-wise comparison) matrix $A=[a_{ij}]$ and a positive vector $w=\left[\begin{array}[c]{ccc}w_{1}&\cdots&w_{n}\end{array}\right]^{T},$ the directed graph $G(A,w)$ has an edge from $i$ to $j$ if and only if $w_{i}\geq a_{ij}w_{j}.$ Vector $w$ is efficient for $A$ if and only if $G(A,w)$ is strongly connected if and only if (new to this literature) $G(A,w)$ contains a Hamiltonian cycle. We show an exact correspondence between such cycles and Hamiltonian cycle products in $A$ that are $\leq 1.$ This facilitates a way to describe all efficient vectors for $A$ via cycle products, and applications, such as that the set of vectors efficient for $A$ determines $A,$ and that, if $G(A,u)$ and $G(A,v)$ present a common cycle, all vectors on the line segment joining $u$ and $v$ are efficient. A sufficient condition for convexity of the efficient set of vectors follows. Order reversals in an efficient vector are also analyzed. The description of the set of efficient vectors for an $n$-by-$n$ column perturbed consistent matrix is given. It is the union of at most $(n-1)(n-2)/2$ convex subsets.
Keywords: cycle products, decision analysis, efficient vector, pair-wise comparisons, reciprocal matrix
MSC2020: 15B48, 05C20, 90B50, 91B06
1 Introduction
An $n$-by-$n$, entry-wise positive matrix $A=[a_{ij}]$ is called reciprocal if $a_{ji}=\frac{1}{a_{ij}}$ for $1\leq i,j\leq n.$ Let $\mathcal{PC}_{n}$ denote the set of all $n$-by-$n$ reciprocal matrices. Such matrices represent independent, pair-wise, ratio comparisons among $n$ alternatives. In a variety of models employing reciprocal matrices $A$, a cardinal ranking vector $w=\left[\begin{array}[c]{cccc}w_{1}&w_{2}&\cdots&w_{n}\end{array}\right]^{T}\in\mathbb{R}_{+}^{n}$ (the entry-wise positive vectors in $\mathbb{R}^{n}$) is to be deduced from $A,$ so that ratios from $w$ approximate the ratio comparisons in $A$ [3, 11, 12, 15, 22, 24, 25, 31].
[…]
There is a graph theoretic method to decide wether $w\in\mathcal{E}(A).$ The graph $G(A,w)$ with vertex set $\{1,\ldots,n\}$ has an edge $i\rightarrow j,$ $1\leq i\neq j\leq n,$ if and only if $w_{i}\geq a_{ij}w_{j}$. First in [6], and then in [17] in a simple matricial way, it was shown that $w\in\mathcal{E}(A)$ if and only if $G(A,w)$ is strongly connected. Since $A$is reciprocal, for every pair of distinct indices $i,j\in\{1,2,\ldots,n\},$ at least one of the edges $i\rightarrow j$ or $j\rightarrow i$ is in $G(A,w).$ A directed graph with at least one of $\{i\rightarrow j,$ $j\rightarrow i\}$ as an edge (and, perhaps both) is called semi-complete [5] (if it never happens that both occur, it is called a tournament [5, 26]); so $G(A,w)$ is always semi-complete. A semi-complete digraph is strongly connected if and only if it contains a full cycle [5], i.e. a Hamiltonian cycle. We add this fact here to the graph theoretic characterization of efficiency.
Theorem 1
Suppose that $A\in\mathcal{PC}_{n}$ and that $w\in\mathbb{R}_{+}^{n}$. The following are equivalent:
  1. (i)
    $w\in\mathcal{E}(A);$
  2. (ii)
    $G(A,w)$ is a strongly connected digraph;
  3. (iii)
    $G(A,w)$ contains a Hamiltonian cycle.
If there is a $w\in\mathbb{R}_{+}^{n}$ such that $A=ww^{(-T)},$ then $A$ is called consistent (otherwise, it is said to be inconsistent). Of course, a consistent matrix is reciprocal and $\mathcal{E}(A)$ just consists of positive multiples of $w.$ Any matrix in $\mathcal{PC}_{2}$ is consistent. Throughout, we focus on the study of $\mathcal{E}(A)$ when $A\in\mathcal{PC}_{n}$ is inconsistent (and, thus, $n\geq 3$).
In [19], a way to generate all vectors in $\mathcal{E}(A)$ (inductively) was given. Here we follow a different approach to characterize $\mathcal{E}(A).$ We connect the Hamiltonian cycles in $G(A,w)$ with the Hamiltonian cycle products $\leq 1$ from matrix $A.$ By a (Hamiltonian) cycle product in $A$ we mean the product of the entries $a_{ij}$ for which $i\rightarrow j$ is an edge of the cycle. There are at most $\frac{(n-1)!}{2}$ of these products $<1$, and exactly this number if no cycle product is $1$. We show that $\mathcal{E}(A)$ is the union of at most $\frac{(n-1)!}{2}$ convex subsets. Each subset is associated with a Hamiltonian cycle product from $A<1$ (when $A$ is inconsistent). We give sufficient conditions for the convexity of $\mathcal{E}(A),$ and for the cone generated by the columns of $A$ (that is, the set of nonzero, nonnegative linear combinations of the columns of $A$) to lie in $\mathcal{E}(A).$ The former implies the latter since any column of $A$ lies in $\mathcal{E}(A)$ [18]. When the latter occurs, the efficiency of the Perron vector of $A$ and of the left singular vector of $A$ (that is, the Perron vector of $AA^{T}$) follows, as these vectors belong to the cone generated by the columns [21].
Using our new description of the efficient vectors for a reciprocal matrix, we show that $A$ is determined by $\mathcal{E}(A)$. Also, we give necessary and sufficient conditions for the existence of an efficient vector associated with a Hamiltonian cycle to have no order reversals with entries along the cycle and show that, in any circumstance, there is always one efficient vector with at most one such order reversal. An efficient vector $w$ for $A\in\mathcal{PC}_{n}$ is said to exhibit an order reversal at $i,j$ if $w_{i}>w_{j}$ ($w_{i}<w_{j}$) when $a_{ij}<1$ ($a_{ij}>1$), $w_{i}=w_{j}$ when $a_{ij}\neq 1$, or $w_{i}\neq w_{j}$ when $a_{ij}=1.$ There is now a considerable literature on order reversal and vectors that minimize the number of order reversals (see, for example, [9, 13, 30]).
If one off-diagonal entry of a consistent matrix, and, correspondingly, its symmetrically placed entry, is changed, so that the resulting $A$ lies in $\mathcal{PC}_{n},$ then $A$ is called a simple perturbed consistent matrix. In prior work [10], $\mathcal{E}(A)$ for any simple perturbed consistent matrix $A\in\mathcal{PC}_{n}$ was determined, and this set is defined by a finite system of linear inequalities on the entries of the vectors, implying that $\mathcal{E}(A)$ is convex for simple perturbed consistent matrices $A$. Since this class of matrices includes $\mathcal{PC}_{3},$ it follows that $\mathcal{E}(A)$ is convex for any $A\in\mathcal{PC}_{3}.$ From the inductive description of the efficient vectors for a reciprocal matrix in [19], it follows that each efficient vector for $A\in\mathcal{PC}_{4}$ is an extension of an efficient vector for one of $3$ $3$-by-$3$ principal submatrices of $A.$ Since these extensions are defined by linear inequalities in the entries of the vector, it follows that the set of efficient vectors for $A\in\mathcal{PC}_{4}$ is the union of at most $3$ convex sets. Later, in [29] the authors illustrated geometrically how $\mathcal{E}(A)$ is the union of $3$ convex sets in the $4$-by-$4$ case.
In [21] we have given examples of matrices obtained from consistent matrices by changing one column (and the corresponding row reciprocally), called column perturbed consistent matrices, for which the set of efficient vectors is not convex. However, it was shown that the cone generated by the columns of such matrices $A$ is contained in $\mathcal{E}(A).$ Note that, for any $n\geq 4,$ examples of matrices $A\in\mathcal{PC}_{n}$ for which the Perron eigenvector is not efficient are known [7, 20]. In these cases, the convex hull of the columns of the matrix is not contained in $\mathcal{E}(A)$, implying that $\mathcal{E}(A)$ is not convex. Here we describe the set of efficient vectors for a column perturbed consistent matrix $A\in\mathcal{PC}_{n}$ and show that it is the union of at most $(n-1)(n-2)/2$ convex sets. The simple perturbed consistent matrices and some type of double perturbed consistent matrices are special cases of column perturbed consistent matrices. As mentioned above, the efficient vectors for the former were studied in [10]. Those for the latter were obtained in [16].
A useful observation is that the set $\mathcal{PC}_{n}$ is closed under both positive diagonal similarity and permutation similarity (that is, monomial similarity). Fortunately, such transformations interface with efficient vectors in a natural way.
Lemma 2
[18] Suppose that $A\in\mathcal{PC}_{n}$ and $w\in\mathcal{E}(A).$ Then, if $D$ is a positive diagonal matrix ($P$ is a permutation matrix), then $Dw\in\mathcal{E}(DAD^{-1})$ ($Pw\in\mathcal{E}(PAP^{T})$).
Note that the graphs $G(A,w)$ and $G(DAD^{-1},Dw)$ coincide.
In the next section, we present a fundamental cycle theorem for efficient vectors of reciprocal matrices. In Section 3, we apply that theorem to show that $A$ may be uniquely reconstructed  from $\mathcal{E}(A)$ and in Section 4 we give some results on order reversals in an efficient vector. Then, in Section 5 we apply the fundamental cycle theorem to study the efficient vectors for a column perturbed consistent matrix.
2 The fundamental cycle theorem for efficient vectors of reciprocal matrices
Here, we present a fundamental theorem, based upon cycles, that explains $\mathcal{E}(A)$ for $A\in\mathcal{PC}_{n}.$ This permits insight into the nature of $\mathcal{E}(A),$ including a sufficient condition for convexity and an application that $A$ is determined by $\mathcal{E}(A).$
2.1 Auxiliary results
Next, we note that, for a fixed $A$, the set of all $w$ such that $G(A,w)$ share a common edge is convex.
Lemma 3
Suppose that $A\in\mathcal{PC}_{n}$and that $u,v\in\mathbb{R}_{+}^{n}$. If $G(A,u)$ and $G(A,v)$ share a common edge, that edge must also occur in $G(A,w)$ for any $w=tu+(1-t)v,$ $t\in(0,1).$
Proof. Suppose that $G(A,u)$ and $G(A,v)$ have a common edge $i\rightarrow j.$ Then $u_{i}\geq a_{ij}u_{j}$ and $v_{i}\geq a_{ij}v_{j},$ so that
$w_{i}$ $=tu_{i}+(1-t)v_{i}\geq ta_{ij}u_{j}+(1-t)a_{ij}v_{j}$
$=a_{ij}(tu_{j}+(1-t)v_{j})=a_{ij}w_{j}.$
As an immediate consequence of Lemma 3 and Theorem 1, we have the following.
Lemma 4
Suppose that $A\in\mathcal{PC}_{n}$ and that $u,v\in\mathbb{R}{}_{+}^{n}$. If $G(A,u)$ and $G(A,v)$ share a common Hamiltonian cycle, then the line segment joining $u$ and $v$ lies in $\mathcal{E}(A).$
The following lemma on Hamiltonian cycles in $A$ along which all entries are $1$ will be helpful.
Lemma 5
Suppose that $A\in\mathcal{PC}_{n}$ is inconsistent and $\mathcal{C}$ is a Hamiltonian cycle in $A.$ If all entries in $A$ along $\mathcal{C}$ are $1,$ then there is a Hamiltonian cycle $\mathcal{C}^{\prime}$ in $A$ such that all entries in $A$ along $\mathcal{C}^{\prime}$ are $\leq 1$ and there is at least one entry $<1.$
Proof. Since a permutation similarity on $A$ keeps the sets of entries in $A$ along the Hamiltonian cycles the same, we may assume, without loss of generality, that $\mathcal{C}$ is the cycle $1\rightarrow 2\rightarrow 3\rightarrow\cdots\rightarrow n-1\rightarrow n\rightarrow 1$ Then, $a_{j,j+1}=a_{j+1,j}=a_{1n}=a_{n1}=1,$ for $j=1,\ldots,n-1.$ Since $A$ is not consistent, it has an entry $<1.$ Let $k$ $(>1)$ be the smallest integer such that either the $k$-th upper-diagonal or the $k$-th lower-diagonal of $A$ has an entry $<1.$ Then, in one of such diagonals, one of the following occurs: i) there is an entry $<1$ followed by an entry $\leq 1$; ii) after some possible entries equal to $1,$ the entries alternate between $<1$ and $>1,$ with the last entry $<1.$ Without loss of generality, we may assume that one of these situations occurs for the $k$-th upper-diagonal (as otherwise we may consider $A^{T}$ instead of $A$). Note that $a_{pq}=1$ for $|p-q|<k.$
Case i) Suppose that $a_{i,i+k}<1$ and $a_{i+1,i+1+k}\leq 1.$ Let
$\mathcal{C}^{\prime}$ $:i\rightarrow i+k\rightarrow i+1\rightarrow i+2\rightarrow\cdots\rightarrow i+k-1\rightarrow i+k+1\rightarrow$
$i+k+2$ $\rightarrow\cdots\rightarrow n\rightarrow 1\rightarrow 2\rightarrow\cdots\rightarrow i-1\rightarrow i\text{.}$
Note that, if $k=2,$ then $a_{i+k-1,i+k+1}=a_{i+1,i+1+k}\leq 1.$
Case ii) We consider two subcases.
Case iia) Suppose that $k>2$ and $a_{n-k,n}<1.$ Let
$\mathcal{C}^{\prime}:n-k\rightarrow n\rightarrow 1\rightarrow 2\rightarrow\cdots\rightarrow n-k-1\rightarrow n-k+1\rightarrow n-k+2\rightarrow\cdots\rightarrow n-1\rightarrow n-k\text{. }$
Case iib) Suppose that $k=2$, $a_{n-2,n}<1$, $a_{n-3,n-1}>1$, etc. Let
$\mathcal{C}^{\prime}$ $:1\rightarrow 2\rightarrow 4\rightarrow\cdots\rightarrow n-2\rightarrow n\rightarrow n-1\rightarrow n-3\rightarrow\cdots\rightarrow 3\rightarrow 1,\text{ if }n\text{ is even,}$,
$\mathcal{C}^{\prime}$ $:1\rightarrow 3\rightarrow 5\rightarrow\cdots\rightarrow n-2\rightarrow n\rightarrow n-1\rightarrow n-3\rightarrow\cdots\rightarrow 4\rightarrow 2\rightarrow 1,\text{ if }n\text{ is odd.}$.
In each case, the cycle $\mathcal{C}^{\prime}$ verifies the claim.
2.2 Description of the set $\mathcal{\varepsilon}_{A}(\mathcal{C})$
In what follows, given $A\in\mathcal{PC}_{n}$ and a Hamiltonian cycle $\mathcal{C}$, we denote by $\mathcal{\varepsilon}_{A}(\mathcal{C})$ the set of vectors $w$ such that $G(A,w)$ contains the cycle $\mathcal{C}$. Note that, by Theorem 1, $\varepsilon_{A}(\mathcal{C})\subseteq\mathcal{E}(A)$ and, by Lemma 4, that $\varepsilon_{A}(\mathcal{C})$ is convex. Next, we describe the sets $\mathcal{\varepsilon}_{A}(\mathcal{C}).$ We denote by $\pi_{1}(A)$ and $\pi_{<1}(A)$ the set of all the Hamiltonian cycles for $A$ with product $\leq 1$ and $<1$, respectively. Of course $\pi_{<1}(A)\subseteq\pi_{1}(A).$
Suppose that $\mathcal{C}$ is the Hamiltonian cycle
$\gamma_{1}\rightarrow\gamma_{2}\rightarrow\gamma_{3}\rightarrow\cdots\rightarrow\gamma_{n}\rightarrow\gamma_{1}.$ (1)
Then,
$a_{\gamma_{n}\gamma_{1}}{\displaystyle\prod\limits_{i=1}^{n-1}}a_{\gamma_{i}\gamma_{i+1}}$
is the cycle product for $\mathcal{C}$ in $A.$ We have that $w\in\mathcal{\varepsilon}_{A}(\mathcal{C})$ if and only if
$\begin{array}[c]{rrr}w_{\gamma_{1}}\geq a_{\gamma_{1}\gamma_{2}}w_{\gamma_{2}},&\quad w_{\gamma_{2}}\geq a_{\gamma_{2}\gamma_{3}}w_{\gamma_{3}},&\\ \ldots,&w_{\gamma_{n-1}}\geq a_{\gamma_{n-1}\gamma_{n}}w_{\gamma_{n}},&\quad w_{\gamma_{n}}\geq a_{\gamma_{n}\gamma_{1}}w_{\gamma_{1}},\end{array}$ (2)
or, equivalently,
$w_{\gamma_{1}}$ $\geq a_{\gamma_{1}\gamma_{2}}w_{\gamma_{2}}\geq\left({\displaystyle\prod\limits_{i=1}^{2}}a_{\gamma_{i}\gamma_{i+1}}\right)w_{\gamma_{3}}\geq$ (3)
$\cdots$ $\geq\left({\displaystyle\prod\limits_{i=1}^{n-1}}a_{\gamma_{i}\gamma_{i+1}}\right)w_{\gamma_{n}}\geq\left(a_{\gamma_{n}\gamma_{1}}{\displaystyle\prod\limits_{i=1}^{n-1}}a_{\gamma_{i}\gamma_{i+1}}\right)w_{\gamma_{1}}.$
For a matricial description of $\varepsilon_{A}(\mathcal{C}),$ consider
$P=\left[\begin{tabular}[c]{c|c}$\begin{array}[c]{c}0\\ \vdots\\ 0\end{array}$&$I_{n-1}$\\ \hline\cr$1$&$\begin{array}[c]{ccc}0&\cdots&0\end{array}$\end{tabular}\ \ \ \ \ \ \ \right]\text{ and }S=\operatorname*{diag}(a_{\gamma_{1}\gamma_{2}},\text{ }a_{\gamma_{2}\gamma_{3}},\text{ }\ldots,\text{ }a_{\gamma_{n-1}\gamma_{n}},\text{ }a_{\gamma_{n}\gamma_{1}}).$
Then (2) is equivalent to $SPw_{\gamma}\leq w_{\gamma},$ in which $w_{\gamma}=\left[\begin{array}[c]{cccc}w_{\gamma_{1}}&w_{\gamma_{2}}&\cdots&w_{\gamma_{n}}\end{array}\right]^{T}.$ Thus $w\in\mathcal{\varepsilon}_{A}(\mathcal{C})$ if and only if $w_{\gamma}$ satisfies the system of linear inequalities
$(I_{n}-SP)w_{\gamma}\geq 0\text{, }w_{\gamma}>0.$
We say that a set of efficient vectors for $A$ is a singleton if it only contains a single (positive) vector (and all its positive multiples).
Lemma 6
Suppose that $A\in\mathcal{PC}_{n}$ and that $\mathcal{C},$ as in (1), lies in $\pi_{1}(A)$. The set of positive solutions $w$ of (2), when any $n-1$ inequalities are taken to be equalities, is a singleton. Moreover, a vector in each of the $n$ singletons is extreme in $\mathcal{\varepsilon}_{A}(\mathcal{C}),$ so that $\mathcal{\varepsilon}_{A}(\mathcal{C})$ is the cone generated by these $n$ vectors.
Proof. Let $k\in\{1,\ldots,n\}.$ Suppose that all inequalities in (2) are taken to be equalities except the $k$-th one. First suppose that $k\neq n.$ Multiplying the equalities, we get
${\displaystyle\prod\limits_{i=1,i\neq k}^{n}}w_{\gamma_{i}}$ $=a_{\gamma_{n}\gamma_{1}}w_{\gamma_{1}}{\displaystyle\prod\limits_{i=1,i\neq k}^{n-1}}\left(a_{\gamma_{i}\gamma_{i+1}}w_{\gamma_{i+1}}\right)$
$\Leftrightarrow w_{\gamma_{k+1}}=\left(a_{\gamma_{n}\gamma_{1}}{\displaystyle\prod\limits_{i=1}^{n-1}}a_{\gamma_{i}\gamma_{i+1}}\right)a_{\gamma_{k+1}\gamma_{k}}w_{\gamma_{k}}.$
Since
$a_{\gamma_{n}\gamma_{1}}{\displaystyle\prod\limits_{i=1}^{n-1}}a_{\gamma_{i}\gamma_{i+1}}\leq 1,$
we get
$w_{\gamma_{k+1}}\leq a_{\gamma_{k+1}\gamma_{k}}w_{\gamma_{k}}\Leftrightarrow w_{\gamma_{k}}\geq a_{\gamma_{k}\gamma_{k+1}}w_{\gamma_{k+1}},$
implying that the $k$-th inequality is satisfied. If $k=n,$ it can be seen in a similar way that the equalities taken from the first $n-1$ inequalities in (2) imply the $n$-th inequality. Clearly, taking one entry $1,$ each $n-1$ equalities in (2) determines a unique positive solution for $w.$ So, we get $n$ (some possibly equal) vectors in $\mathcal{\varepsilon}_{A}(\mathcal{C})$ that are precisely the extreme points of the set.
Next we show an important fact about the structure of $\mathcal{\varepsilon}_{A}(\mathcal{C})$.
Lemma 7
Suppose that $A\in\mathcal{PC}_{n}$ and that $\mathcal{C}\in\pi_{1}(A)$. Then $\mathcal{\varepsilon}_{A}(\mathcal{C})$ is a singleton if and only if the cycle product for $\mathcal{C}$ in $A$ is $1$ Moreover, if $A$ is inconsistent and the product for $\mathcal{C}$ in $A$ is $1,$ then $\varepsilon_{A}(\mathcal{C})\subseteq\varepsilon_{A}(\mathcal{C}^{\prime}),$ for some $\mathcal{C}^{\prime}\in\pi_{<1}(A).$
Proof. Since a diagonal similarity does not change any cycle product in $A$, and by a diagonal similarity on $A$ any $n-1$ entries along $\mathcal{C}$ may be made $1$ (and the remaining entry is $\leq 1,$ since $\mathcal{C}\in\pi_{1}(A)$), taking into account Lemma 2, we assume that this situation occurs in order to prove the result. When the product for $\mathcal{C}$ in $A$ is $1,$ all entries in $A$ along $\mathcal{C}$ are $1$ Then, from (2) (with $\mathcal{C}$ as in (1)), $w\in\varepsilon_{A}(\mathcal{C})$ if and only, up to a positive factor, $w$ is the vector of $1$’s, implying that $\mathcal{\varepsilon}_{A}(\mathcal{C})$ is a singleton. When the product of $\mathcal{C}$ in $A$ is $<$ $1,$ by Lemma 6, there is a $w$ satisfying (2) in which the first $n-1$ inequalities are taken to be equalities. Then, the $n$-th inequality is strict. Any vector $w^{\prime}$ obtained from $w$ by a sufficiently small decrease of $w_{\gamma_{n}},$ and agreeing with $w$ in all other components, still satisfies (2) (with $w_{\gamma_{n}}^{\prime}$ instead of $w_{\gamma_{n}}$). So $\mathcal{\varepsilon}_{A}(\mathcal{C})$ is not a singleton. This completes the proof of the first claim.
Suppose that the product along $\mathcal{C}$ in $A$ (inconsistent) is $1,$ in which case all vectors $w$ in $\varepsilon_{A}(\mathcal{C})$ are constant. By Lemma 5, there is a Hamiltonian cycle $\mathcal{C}^{\prime}$ in $A$ such that all entries in $A$ along $\mathcal{C}^{\prime}$ are $\leq 1$ and there is at least one entry $<1.$ For $\mathcal{C}^{\prime}$ as in (1), it follows that $w$ satisfies (2), implying that $w\in\varepsilon_{A}(\mathcal{C}^{\prime}).$ When $\mathcal{C}^{\prime}\in\pi_{<1}(A),$ the last claim follows.
We observe that, if the product along a Hamiltonian cycle $\mathcal{C}$ in $A$ is $1,$ then so is the product along $\mathcal{C}^{\prime},$ and $\mathcal{\varepsilon}_{A}(\mathcal{C})=\mathcal{\varepsilon}_{A}(\mathcal{C}^{\prime}),$ in which $\mathcal{C}^{\prime}$ is the reverse cycle of $\mathcal{C}$.
2.3 Main result
Now, we may give the fundamental theorem on the cycle structure of reciprocal matrices and their efficient vectors. If $A\in\mathcal{PC}_{n}$ is consistent, then $\mathcal{E}(A)$ is the singleton generated by any column of $A,$ otherwise, since each column is in $\mathcal{E}(A)$ and not all columns are proportional, $\mathcal{E}(A)$ is not a singleton. For simplicity, we assume that $A$ is inconsistent.
Theorem 8
Suppose that $A\in\mathcal{PC}_{n}$ is inconsistent. If $w\in\mathcal{E}(A)$ and $\mathcal{C}$ is a Hamiltonian cycle in $G(A,w),$ then $\mathcal{C}\in\pi_{1}(A).$ On the other hand, if $\mathcal{C}\in\pi_{1}(A),$ then the set
$\mathcal{\varepsilon}_{A}(\mathcal{C})=\left\{w:G(A,w)\text{ contains }\mathcal{C}\right\}$
is a nonempty, convex subset of $\mathcal{E}(A).$ Moreover,
$\mathcal{E}(A)={\displaystyle\bigcup\limits_{C\in\pi_{<1}(A)}}\mathcal{\varepsilon}_{A}(\mathcal{C}).$ (4)
Proof. If $w\in\mathcal{E}(A)$ and $\mathcal{C}$, as in (1), is a Hamiltonian cycle in $G(A,w),$ then $w\in\varepsilon_{A}(\mathcal{C}).$Hence, (3) holds and the inequality between the left most and the right most expressions imply
$a_{\gamma_{n}\gamma_{1}}{\displaystyle\prod\limits_{i=1}^{n-1}}a_{\gamma_{i}\gamma_{i+1}}\leq 1,$
which means that $\mathcal{C}\in\pi_{1}(A).$
Suppose that $\mathcal{C}\in\pi_{1}(A).$ The set $\varepsilon_{A}(\mathcal{C})$ is nonempty by Lemma 6. As, for $w\in\mathcal{\varepsilon}_{A}(\mathcal{C})$, $G(A,w)$ contains the cycle $\mathcal{C}$, by Theorem 1, $w\in\mathcal{E}(A).$ Also, note that, as all such $G(A,w)$ contain $\mathcal{C}$, the set is convex by Lemma 4 (alternatively, (polyhedral) convexity follows from the fact that $w\in\mathcal{\varepsilon}_{A}(\mathcal{C})$ is determined by the finite system of linear inequalities (2)). Since, by Theorem 1, each $w\in\mathcal{E}(A)$lies in $\varepsilon_{A}(\mathcal{C})$ for some $\mathcal{C}$, and, by the first part of the proof, $\mathcal{C}\in\pi_{1}(A)$, it follows that
$\mathcal{E}(A)={\displaystyle\bigcup\limits_{C\in\pi_{1}(A)}}\mathcal{\varepsilon}_{A}(\mathcal{C}).$ (5)
Now (4) follows taking into account the last claim in Lemma 7. This conclusion also follows from the following alternate argument. Since $\mathcal{E}(A)$ is connected [6] and is not a singleton (as $A$ is inconsistent), it follows that any singleton $\mathcal{\varepsilon}_{A}(\mathcal{C})$ in the union in (5) should be contained in a non-singleton $\mathcal{\varepsilon}_{A}(\mathcal{C}^{\prime}),$ which, by Lemma 7, implies that $\mathcal{C}^{\prime}\in\pi_{<1}(A).$
Note that $\pi_{<1}(A)$ has no more than $\frac{(n-1)!}{2}$ elements and has less if there are Hamiltonian cycles from $A$ with product $1$. When $n=3,$ then $\frac{(n-1)!}{2}=1$ which means that $\mathcal{E}(A)=\varepsilon_{A}(\mathcal{C})$ for some cycle $\mathcal{C}\in$ $\pi_{1}(A),$ which is another way to see that $\mathcal{E}(A)$ is convex for $A\in\mathcal{PC}_{3}.$
We next give some consequences of Theorem 8. Since the results can be trivially verified for consistent matrices, we state them for general reciprocal matrices.
Corollary 9
Suppose $A\in\mathcal{PC}_{n}.$ If there exists a Hamiltonian cycle that lies in $G(A,w)$ for every $w\in\mathcal{E}(A),$ then $\mathcal{E}(A)$ is convex. In particular, this holds if there is just one Hamiltonian cycle product $<1$ in $A$.
Observe that the first claim in Corollary 9 is also a consequence of Lemma 4.
Corollary 10
Suppose $A\in\mathcal{PC}_{n}.$ If all columns of $A$ are in $\mathcal{\varepsilon}_{A}(\mathcal{C})$ for some $\mathcal{C}\in\pi_{1}(A)$, then the cone generated by the columns of $A$ is contained in $\mathcal{E}(A)$. In particular, the Perron vector and the singular vector of $A$ are efficient for $A.$
It is known that in general $\mathcal{E}(A)$ is not closed under component-wise geometric mean [18]. However, we note that each subset $\mathcal{\varepsilon}_{A}(\mathcal{C})$ is, because it is defined by inequalities (2).
Corollary 11
Suppose $A\in\mathcal{PC}_{n}.$ If there exists a cycle that lies in $G(A,w)$ for every $w\in\mathcal{E}(A),$ then $\mathcal{E}(A)$ is closed under component-wise geometric mean.
We give next an example illustrating the previous results.
Example 12
Let
$A=\left[\begin{array}[c]{cccc}1&2&1&\frac{1}{2}\\ \frac{1}{2}&1&2&1\\ 1&\frac{1}{2}&1&2\\ 2&1&\frac{1}{2}&1\end{array}\right].$
Note that there are $2$ cycle products in $A$ equal to $1,$
$\mathcal{C}_{1}$ $:1\rightarrow 2\rightarrow 4\rightarrow 3\rightarrow 1,$
$\mathcal{C}_{2}$ $:1\rightarrow 3\rightarrow 2\rightarrow 4\rightarrow 1,$
and their reverses, and the cycle product $<1$
$\mathcal{C}_{3}:1\rightarrow 4\rightarrow 3\rightarrow 2\rightarrow 1.$
By Theorem 8, $\mathcal{E}(A)=\mathcal{\varepsilon}_{A}(\mathcal{C}_{3}),$ and, so, $\mathcal{E}(A)$ is convex. In particular, the Perron vector and the singular vector of $A$ are efficient for $A.$ Using (2), we can see that
$\mathcal{E}(A)=\mathcal{\varepsilon}_{A}(\mathcal{C}_{3})=\left\{w\in\mathbb{R}_{+}^{4}:w_{1}\geq\frac{1}{2}w_{4}\geq\frac{1}{4}w_{3}\geq\frac{1}{8}w_{2}\geq\frac{1}{16}w_{1}\right\}.$
We also have
$\mathcal{\varepsilon}_{A}(\mathcal{C}_{1})$ =$=\left\{w\in\mathbb{R}_{+}^{4}:w_{1}=2w_{2}=2w_{4}=w_{3}\right\},$
$\mathcal{\varepsilon}_{A}(\mathcal{C}_{2})$ =$=\left\{w\in\mathbb{R}_{+}^{4}:w_{1}=w_{3}=\frac{1}{2}w_{2}=\frac{1}{2}w_{4}\right\},$
which can be verified to be contained in $\mathcal{\varepsilon}_{A}(\mathcal{C}_{3}),$ as expected. We observe that $A$ is not a simple perturbed consistent matrix, as it is nonsingular. Thus, this example illustrates that the convexity of $\mathcal{E}(A)$ may occur for other matrices $A$ than the simple perturbed consistent matrices.
3 $\mathcal{E}(A)$ uniquely determines $A$
Next we turn our attention to an interesting application of the fundamental theorem, other than the description of all efficient vectors.
Of course, $A\in\mathcal{PC}_{n}$ determines its efficient set $\mathcal{E}(A).$ We show that the converse is also so: $A\in\mathcal{PC}_{n}$ is uniquely determined by $\mathcal{E}(A).$ Our strategy is to show that, for every pair of distinct indices $i,j\in\{1,2,\ldots,n\},$ there is a vector $w\in\mathcal{E}(A)$ such that both $i\rightarrow j$ and $j\rightarrow i$ are edges in $G(A,w)$. This, means $w_{i}\geq a_{ij}w_{j}$ and $w_{j}\geq a_{ji}w_{i}.$ Since $a_{ij}a_{ji}=1,$ we must have $w_{i}=a_{ij}w_{j}$, i.e. $a_{ij}=\frac{w_{i}}{w_{j}},$ so that $a_{ij}$ (and, thus, $a_{ji}$) are determined by $w\in\mathcal{E}(A).$
Lemma 13
If $A\in\mathcal{PC}_{n}$ and $i,j\in\{1,\ldots,n\}$ with $i\neq j$, then there is a $\mathcal{C}\in\pi_{1}(A)$ that uses either the edge $i\rightarrow j$ or the edge $j\rightarrow i.$
Proof. Pick any cycle using $i\rightarrow j.$ If its product is $\leq 1,$ we are done. If not, the reverse of this cycle has the reciprocal product that is $\leq 1$ and uses the edge $j\rightarrow i.$
Lemma 14
Suppose that $A\in\mathcal{PC}_{n}$, that $w\in\mathbb{R}_{+}^{n}$ and that $i,j\in\{1,\ldots,n\}$ with $i\neq j$. If $\mathcal{C}$ is a Hamiltonian cycle in $G(A,w)$ that contains the edge $i\rightarrow j,$ then there is a vector $w^{\prime}\in\mathbb{R}_{+}^{n}$ such that $\mathcal{C}$ lies in $G(A,w^{\prime})$ and such that $G(A,w^{\prime})$ includes the additional edge $j\rightarrow i.$
Proof. Since $i\rightarrow j$ is an edge of $G(A,w),$ $w_{i}\geq a_{ij}w_{j}.$ Now define $w^{\prime}$ to agree with $w$ in all components except the $j$-th, and define $w_{j}^{\prime}=\frac{w_{i}}{a_{ij}}.$ Since $w_{j}^{\prime}\geq w_{j},$ $\mathcal{C}$ also lies in $G(A,w^{\prime}).$ Since $w^{\prime}$ satisfies $w_{i}^{\prime}=a_{ij}w_{j}^{\prime},$ we have $w_{j}^{\prime}=a_{ji}w_{i}^{\prime},$ verifying that $j\rightarrow i$ is also an edge of $G(A,w^{\prime}).$
Now we may apply our strategy to verify the following.
Theorem 15
Let $A,B\in\mathcal{PC}_{n}$. If $\mathcal{E}(A)=\mathcal{E}(B)$ then $A=B.$
Proof. If $\mathcal{E}(A)$ is a singleton, then $A$ is consistent and the columns of $A$ are proportional to any vector in $\mathcal{E}(A)$ and are uniquely determined since all entries on the main diagonal of $A$ are $1$ Suppose that $\mathcal{E}(A)$ is not a singleton ($A$ is inconsistent) and $i,j\in\{1,2,\ldots,n\}$ with $i\neq j.$ By Lemma 13 and Theorem 8, there is a Hamiltonian cycle $\mathcal{C}$ containing one of the edges $i\rightarrow j$ or $j\rightarrow i$ and a vector $w\in\mathcal{E}(A)$ such that $G(A,w)$ contains $\mathcal{C}$. Then, Lemma 14 and Theorem 1 guarantee a $w^{\prime}\in\mathcal{E}(A)$ with $w_{i}^{\prime}=a_{ij}w_{j}^{\prime}$, implying that $a_{ij}=\frac{w_{i}^{\prime}}{w_{j}^{\prime}}.$ Thus, the $i,j$ entry of $A$ is determined. Since the pair $i,j$ is arbitrary, $A$ is fully determined.
4 Order reversals
Here we give some consequences of Theorem 8 regarding the existence of no order reversals in an efficient vector. Of course, if $A$ is consistent, no efficient vector for $A$ exhibits an order reversal, as $\frac{w_{i}}{w_{j}}=a_{ij}$ for all $i,j.$
Theorem 16
Suppose that $A\in\mathcal{PC}_{n}$ is inconsistent and $\mathcal{C}\in\pi_{1}(A).$ Then, there is a $w\in\varepsilon_{A}(\mathcal{C)}$ that exhibits no order reversals with entries of $A$ along $\mathcal{C}$ if and only if either there is an entry of $A$ $>1$ along $\mathcal{C}$ or the cycle product is $1$ Otherwise, there is a $w\in\varepsilon_{A}(\mathcal{C)}$ with at most $1$ order reversal along $\mathcal{C}$.
Proof. Suppose that there is a $w\in\varepsilon_{A}(\mathcal{C)}$ that exhibits no order reversals with entries of $A$ along $\mathcal{C}$ and all entries along $\mathcal{C}$ are $\leq 1,$ with at least one inequality strict. Then $\frac{w_{i}}{w_{j}}$ would be $\leq 1$ for the efficient vector $w$ and all $i,j$ along the cycle, with at least one inequality strict, implying
$1={\displaystyle\prod\limits_{i,j\text{ along }\mathcal{C}}}\frac{w_{i}}{w_{j}}<1,$
a contradiction. So, an entry $>1$ or all entries $1$ along the cycle is necessary.
Suppose there is an entry $>1$ or all entries are $1$ along $\mathcal{C}$. For $\mathcal{C}$ as in (1), suppose, without loss of generality, that $a_{\gamma_{n}\gamma_{1}}\geq 1$, with $a_{\gamma_{n}\gamma_{1}}=1$ if and only if all entries along the cycle are $1$ Now, use the inequalities (2) and take the first $n-1$ of them to be equalities, so as to define $w,$ taking $w_{1}=1$ (Lemma 6). The last inequality (which is implied by the first $n-1$ inequalities) then insures that there is no order reversal at $i_{n},i_{1}$. In fact, $a_{\gamma_{n}\gamma_{1}}>1$ implies $\frac{w_{\gamma_{n}}}{w_{\gamma_{1}}}>1$. Also, $a_{\gamma_{n}\gamma_{1}}=1$ implies, by hypothesis, all entries of $A$ along the cycle equal to $1,$ and thus, by the $n-1$ equalities, $\frac{w_{\gamma_{j+1}}}{w_{\gamma_{j}}}=1,$ $11\leq j\leq n-$ 1 . Then, $\frac{w_{\gamma_{n}}}{w_{\gamma_{1}}}=1.$ There is no order reversal elsewhere along the cycle because of the equalities. The same construction insures at most one order reversal without the assumption of an entry $>1$ or all entries $1$ along $\mathcal{C}.$
5 Efficient vectors for column perturbed consistent matrices
In this section, we study the efficient vectors for column perturbed consistent matrices [20, 21]. Based on Lemma 2 and the following observation, we may assume that these matrices have a simple form. By $J_{k}$ we denote the $k$-by-$k$ matrix with all entries $1$
Lemma 17
If $B\in\mathcal{PC}_{n}$ is a column perturbed consistent matrix, then $B$ is monomially similar to a matrix of the form
$A=\left[\begin{array}[c]{cc}1&\begin{array}[c]{cccc}a_{12}&a_{13}&\cdots&a_{1n}\end{array}\\ \begin{array}[c]{c}\frac{1}{a_{12}}\\ \frac{1}{a_{13}}\\ \vdots\\ \frac{1}{a_{1n}}\end{array}&J_{n-1}\end{array}\right]\in\mathcal{PC}_{n}.$ (6)
Proof. Suppose that $B$is obtained from the consistent matrix $ww^{(-T)}$ by modifying, say, row and column $i.$ Then, for $D^{-1}=\operatorname*{diag}(w),$ we have that $DBD^{-1}$ has all entries equal $1$ except those in row and column $i.$ Then $DBD^{-1}$ is permutationally similar to a reciprocal matrix with all entries $1$ except those in the first row and column.
Before we give the description of the efficient vectors for a column perturbed consistent matrix, we illustrate it with an example.
Example 18
Let
$A=\left[\begin{array}[c]{ccccc}1&\frac{1}{5}&\frac{1}{4}&2&3\\ 5&1&1&1&1\\ 4&1&1&1&1\\ \frac{1}{2}&1&1&1&1\\ \frac{1}{3}&1&1&1&1\end{array}\right].$
We have no cycle products equal to $1,$ since $a_{1i}a_{j1}\neq 1$ for $i\neq j.$ So, we have exactly $12$ cycle products $<1.$ They can be written as
$\mathcal{C}_{\gamma}:1\rightarrow\gamma_{2}\rightarrow\gamma_{3}\rightarrow\gamma_{4}\rightarrow\gamma_{5}\rightarrow 1,$
with $a_{1\gamma_{2}}a_{\gamma_{5}1}<1.$ Then
$\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma})=\left\{w:w_{1}\geq a_{1\gamma_{2}}w_{\gamma_{2}}\wedge w_{\gamma_{2}}\geq w_{\gamma_{3}}\geq w_{\gamma_{4}}\geq w_{\gamma_{5}}\wedge w_{\gamma_{5}}\geq a_{\gamma_{5}1}w_{1}\right\}.$
We have that $\gamma$ can be one of the following permutations of $\{2,3,4,5\}$:
$\gamma^{(1)}$ $:2453,\quad\gamma^{(2)}:2543,\quad\gamma^{(3)}:2354,\quad\gamma^{(4)}:2534,\quad\gamma^{(5)}:2345,\quad\gamma^{(6)}:2435,\quad$
$\gamma^{(7)}$ $:3254,\quad\gamma^{(8)}:3524,\quad\gamma^{(9)}:3245,\quad\gamma^{(10)}:3425,\quad\gamma^{(11)}:4235,\quad\gamma^{(12)}:4325.$
Then,
$\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(1)}})\cup\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(2)}})$ =$=\left\{w\in\mathbb{R}_{+}^{5}:a_{21}w_{1}\geq w_{2}\geq w_{4},w_{5}\geq w_{3}\geq a_{31}w_{1}\right\},$
$\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(3)}})\cup\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(4)}})$ =$=\left\{w\in\mathbb{R}_{+}^{5}:a_{21}w_{1}\geq w_{2}\geq w_{3},w_{5}\geq w_{4}\geq a_{41}w_{1}\right\},$
$\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(5)}})\cup\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(6)}})$ =$=\left\{w\in\mathbb{R}_{+}^{5}:a_{21}w_{1}\geq w_{2}\geq w_{3},w_{4}\geq w_{5}\geq a_{51}w_{1}\right\},$
$\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(7)}})\cup\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(8)}})$ =$=\left\{w\in\mathbb{R}_{+}^{5}:a_{31}w_{1}\geq w_{3}\geq w_{2},w_{5}\geq w_{4}\geq a_{41}w_{1}\right\},$
$\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(9)}})\cup\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(10)}})$ =$=\left\{w\in\mathbb{R}_{+}^{5}:a_{31}w_{1}\geq w_{3}\geq w_{2},w_{4}\geq w_{5}\geq a_{51}w_{1}\right\},$
$\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(11)}})\cup\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(12)}})$ =$=\left\{w\in\mathbb{R}_{+}^{5}:a_{41}w_{1}\geq w_{4}\geq w_{2},w_{3}\geq w_{5}\geq a_{51}w_{1}\right\}.$
By Theorem 8,
$\mathcal{E}(A)={\displaystyle\bigcup\limits_{i=1}^{6}}\left(\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(2i-1)}})\cup\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(2i)}})\right).$
Note that each set $\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(2i-1)}})\cup\mathcal{\varepsilon}_{A}(\mathcal{C}_{\gamma^{(2i)}})$ is convex. So, in this case $\mathcal{E}(A)$ is the union of $6$ convex sets.
For $A\in\mathcal{PC}_{n}$ and $i,j\in\{1,\ldots,n\},$ $i\neq j,$ let
$\mathcal{\varepsilon}_{ij}(A)=\left\{w\in\mathbb{R}_{+}^{n}:a_{i1}w_{1}\geq w_{i}\geq w_{k}\geq w_{j}\geq a_{j1}w_{1}\text{, }k\neq 1,i,j\right\}.$
Since $w$ in $\mathcal{\varepsilon}_{ij}(A)$ is defined by a finite number of linear inequalities in its (positive) entries, we have the following.
Lemma 19
For $A\in\mathcal{PC}_{n}$ and $i,j\in\{1,\ldots,n\},$ $i\neq j,$ $\mathcal{\varepsilon}_{ij}(A)$ is convex.
Lemma 20
Let $i,j\in\{2,\ldots,n\},$ $i\neq j.$ If $A\in\mathcal{PC}_{n}$ is as in (6) with $a_{1i}a_{j1}<1,$ then $\mathcal{\varepsilon}_{ij}(A)\subseteq\mathcal{E}(A).$ In particular, $\mathcal{\varepsilon}_{ij}(A)$ is the subset of $\mathcal{E}(A)$ that comes from the cycles of $\pi_{<1}(A)$ of the form $1\rightarrow i\rightarrow\cdots\rightarrow j\rightarrow 1$
Proof. Denote by $\mathcal{C}_{ij}$ the set of all Hamiltonian cycles $\mathcal{C}$ of the form $1\rightarrow i\rightarrow\gamma_{3}\rightarrow\cdots\rightarrow\gamma_{n-1}\rightarrow j\rightarrow 1$ (there are $(n-3)!$). Suppose that $a_{1i}a_{j1}<1,$ so that the product for any cycle $\mathcal{C}$ in $\mathcal{C}_{ij}$ is $<1.$ We have
$\mathcal{\varepsilon}_{A}(\mathcal{C})=\left\{w\in\mathbb{R}_{+}^{n}:a_{i1}w_{1}\geq w_{i}\geq w_{\gamma_{3}}\geq\cdots\geq w_{\gamma_{n-1}}\geq w_{j}\geq a_{j1}w_{1}\right\}.$
Then
$\mathcal{\varepsilon}_{ij}(A)={\displaystyle\bigcup\limits_{\mathcal{C\in C}_{ij}}^{n}}\mathcal{\varepsilon}_{A}(\mathcal{C}).$
Since, by Theorem 8, $\mathcal{\varepsilon}_{A}(\mathcal{C})\subseteq\mathcal{E}(A),$ for each $\mathcal{C\in C}_{ij},$ the claim follows.
We give next the main result of this section.
Theorem 21
If $A\in\mathcal{PC}_{n}$ is inconsistent of the form (6), then
$\mathcal{E}(A)={\displaystyle\bigcup\limits_{(i,j)\in\mathcal{N}}}\mathcal{\varepsilon}_{ij}(A),$ (7)
for $\mathcal{N}=\{(i,j):i,j\in\{2,\ldots,n\}$ and $a_{1i}a_{j1}<1\}.$
Proof. Let $\mathcal{C}$ be a Hamiltonian cycle in $A.$ By Lemma 20, $\mathcal{\varepsilon}_{ij}(A),$ with $(i,j)\in\mathcal{N}$, is contained in $\mathcal{E}(A),$ proving the inclusion $\supseteq$ in (7). On the other hand, by Theorem 8, if $w\in\mathcal{E}(A),$ then $w\in\varepsilon_{A}(\mathcal{C})$ for some $\mathcal{C}\in\pi_{<1}(A).$ Then, for $\mathcal{C}:1\rightarrow i\rightarrow\gamma_{3}\rightarrow\cdots\rightarrow\gamma_{n-1}\rightarrow j\rightarrow 1$, we have $a_{1i}a_{j1}<1.$ Since the cycle product for $\mathcal{C}$ in $A$ is $a_{1i}a_{j1},$ by Lemma 20, $\varepsilon_{A}(\mathcal{C})\subseteq\mathcal{\varepsilon}_{ij}(A)$.
The set $\mathcal{N}$ has at most $\frac{(n-1)(n-2)}{2}$ elements and has exactly this number if $a_{12},\ldots,a_{1n}$ are pairwise distinct.
Corollary 22
If $B\in\mathcal{PC}_{n}$ is a general inconsistent column perturbed consistent matrix, then $\mathcal{E}(B)$ is the union of (at most) $\frac{(n-1)(n-2)}{2}$ convex sets.
Proof. By previous observations, there is an $n$-by-$n$ monomial matrix $S$ such that $A=S^{-1}BS$ is as in (6). By Lemma 2, $\mathcal{E}(B)=S\mathcal{E}(A).$ Thus, by Theorem 21,
$\mathcal{E}(B)={\displaystyle\bigcup\limits_{(i,j)\in\mathcal{N}}}S\mathcal{\varepsilon}_{ij}(A),$
with $\mathcal{N}$ as in Theorem 21. Since, by Lemma 19, $\mathcal{\varepsilon}_{ij}(A)$ is convex then $S\mathcal{\varepsilon}_{ij}(A)$ is convex [21].
We finally show that the description of the efficient vectors for a simple perturbed consistent matrix given in [10] is a simple consequence of the results developed in this paper. Suppose that $A\in\mathcal{PC}_{n},$ $n>2,$ is as in (6), with $a_{12}=\cdots=a_{1,n-1}=1$ and $a_{1n}>1$ (if $a_{1n}=1,$ $A$ is consistent), which can be assumed for the purpose of studying the efficient vectors for a simple perturbed consistent matrix, by Lemma 2. Then, the Hamiltonian cycles whose products from $A$ are $<1$ are of the form $\mathcal{C}:1\rightarrow\ell\rightarrow\gamma_{3}\rightarrow\cdots\rightarrow\gamma_{n-1}\rightarrow n\rightarrow 1$, $\ell=2,\ldots,n-1.$ We have
$\mathcal{\varepsilon}_{\ell n}(A)=\left\{w\in\mathbb{R}_{+}^{n}:w_{1}\geq w_{\ell}\geq w_{k}\geq w_{n}\geq a_{n1}w_{1}\text{, }k\neq 1,\ell,n\right\}.$
Then,
${\displaystyle\bigcup\limits_{\ell=2}^{n-1}}\mathcal{\varepsilon}_{\ell n}(A)=\left\{w\in\mathbb{R}_{+}^{n}:w_{1}\geq w_{k}\geq w_{n}\geq a_{n1}w_{1}\text{,}k\neq 1,n\right\}.$ (8)
By Theorem 21, $\mathcal{E}(A)$ is the set (8), as claimed in [10].